using the distance formula
how would u find a in the following problem
(-4,1)(a,8);d=square root of 50

Answer by checkley77(12569)

X DISTANCE=A+4
Y DISTANCE=8-1=7
(A+4)^2+7^2=SQRT50^2
A^2+8A+16+49=50
A^2+8X+65-50=0
A^2+8X+15=0
(A+5)(A+3)=0
A+5=0
A=-5 ANSWER
A+3=0
A=-3 ANSWWER